OpenStudy (anonymous):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of (f + g)(3)?

OpenStudy (anonymous):

the answers are −4 4 −32 8

OpenStudy (anonymous):

f(x) = x2 − 13 do you mean f(x) = x^2 − 13

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

if so then (x^2 − 13) + ( x + 5) x^2+x-8 add like terms 3^2+3-8 plug in 3 where there are xs =4

OpenStudy (anonymous):

can you help me with this one : Consider the functions f(x) = 13x2 and g(x) = 7x − 5. What is the value of (f − g)(3)?

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

subtract the 2 functions and plug in 3 13*x^2-7*x+5 101

OpenStudy (anonymous):

I know this is alot but could you help me with this one to ? Consider the functions f(x) = 8x + 2 and g(x) = 5x − 3. What is the value of ( )(4)? Please show all of your work for full credit.

OpenStudy (anonymous):

of ( )(4) ^ ?

OpenStudy (anonymous):

f[g(−4)]?

OpenStudy (anonymous):

f[g(−4)] if this is right and you sure it is that then ooh you plug in the whole 5x- 3 into 8x+2 in the x so 8*(5x- 3)+2 40*x-22 distribute the 8 into 5x and - 3 then add the +2 Plug in - 4 -182

OpenStudy (anonymous):

woops, its f/g @timo86m

OpenStudy (anonymous):

f(x) = 8x + 2 and g(x) = 5x − 3 (-4) (8x+2) / (5x-3) just divide then (8(-4)+2)/(5(-4)-3) plug in the -4 where the xs are at 30/23